OpenStudy (anonymous):

Medals! I just need an answer check. A 20-ft pole casts a shadow 12 ft long. How tall is a nearby building that casts a shadow 20 ft long?

OpenStudy (whpalmer4):

Okay, what's your answer?

OpenStudy (anonymous):

Whoops, forgot to put it... My bad. 33.3 ft. tall

OpenStudy (anonymous):

Sorry about that.

OpenStudy (whpalmer4):



OpenStudy (anonymous):

I see that I am not correct.

OpenStudy (whpalmer4):

Yes, by similar triangles that should be correct: \[\frac{20}{12} = \frac{x}{20}\]\[20*20=12x\]\[x=\frac{100}3\approx 33.3\]

OpenStudy (anonymous):

Wait, so my answer is correct?

OpenStudy (whpalmer4):

Well, I think so. Assuming the sun is far enough away that there's no parallax error anything like that, the angle between the height and the hypotenuse should be identical in both figures and we should have similar triangles so the ratio should be the same for corresponding sides

OpenStudy (anonymous):

Okay, thank you very much! :)